OpenStudy (anonymous):

1) The roof lines of a building can be described by the system of equations below, where the floor is represented by the x-axis and the y-axis and the height of the building is along the z-axis. Measurements are in feet. Find the point of intersection of the roof lines, using an ordered triple in the form (x, y, z). x+y+z=53 3x-2y+2=69 -x+2y-7=-59 2) A support column will be placed under the intersection point. The column must be ____ feet to reach from the floor to the intersection point. (Enter only the number.)

OpenStudy (anonymous):

hi can you help me?

OpenStudy (anonymous):

would i have to do elimination first?

OpenStudy (anonymous):

x+y+z=53 3x-2y+2z=69 -x+2y-7z=-59 This right?

OpenStudy (anonymous):

idk im lost

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

last one is a z not 7z

OpenStudy (anonymous):

last one is a z not 7z

OpenStudy (anonymous):

would it be i solver for x first?

OpenStudy (anonymous):

x+y+z=53 3x-2y+2z=69 -x+2y+z=-59 z=-x-y+53

OpenStudy (anonymous):

3x-2y+2(-x-y+53)=69 -x+2y-x-y+53=-59

OpenStudy (anonymous):

last one is a z not 7z

OpenStudy (anonymous):

what would i need to do solve for a variable first?

OpenStudy (anonymous):

x= 475/7?

OpenStudy (anonymous):

I got x=485/7

OpenStudy (anonymous):

i got x= 475/7 and y= 166/7

OpenStudy (anonymous):

Hold on, let me work it out

OpenStudy (anonymous):

i could of got it wrong

OpenStudy (anonymous):

then got z=-8.86

OpenStudy (anonymous):

\[\array{x + y + z == 53 \\ 3 x - 2 y + 2 z = 69 \\ -x + 2 y + z == -59}\] z = -x - y + 53 3 x - 2 y + 2 (-x - y + 53) = 69 -x + 2 y + (-x - y + 53) =-59 \[\array{x-4y\text{==}-37 \\-2x+y\text{==}-112}\] \[\array{2x - 8 y = -74\\ -2 x + y = -112}\] -7y=-186 y=186/7

OpenStudy (anonymous):

oh i messed up then

OpenStudy (anonymous):

whats the awnser for x and z?

OpenStudy (anonymous):

question from 8 months ago?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what's the answer to part 2 of this problem???? please im struggling so bad on this

OpenStudy (anonymous):

Im on the same one.. I got the first part but like you guys the second part makes no sense. Ugh :/